### v2

# From SGD to Spectra: A Theory of Neural Network Weight Dynamics

Abstract
Deep neural networks have revolutionized machine learning, yet their training dynamics remain theoretically unclear—we develop a continuous-time, matrix-valued stochastic differential equation (SDE) framework that rigorously connects the microscopic dynamics of SGD to the macroscopic evolution of singular-value spectra in weight matrices. We derive exact SDEs showing that squared singular values follow Dyson Brownian motion with eigenvalue repulsion, and characterize stationary distributions as gamma-type densities with power-law tails, providing the first theoretical explanation for the empirically observed ‘bulk+tail’ spectral structure in trained networks. Through controlled experiments on transformer and MLP architectures, we validate our theoretical predictions and demonstrate quantitative agreement between SDE-based forecasts and observed spectral evolution, providing a rigorous foundation for understanding why deep learning works.
Machine Learning, ICML
1 Introduction
Deep neural networks have fundamentally transformed machine learning, achieving unprecedented performance across diverse domains (Krizhevsky et al., 2012; Vaswani et al., 2017; Jumper et al., 2021). Yet despite their empirical success, our theoretical understanding of how neural networks learn remains remarkably incomplete (Zhang et al., 2017; Neyshabur et al., 2017). Central to this understanding is the evolution of weight matrices, whose spectral properties—the distribution and dynamics of singular values—provide deep insights into optimization dynamics, generalization behavior, and implicit regularization (Pennington et al., 2017; Martin and Mahoney, 2018).
At initialization, weight matrices exhibit well-characterized random matrix statistics described by the Marchenko-Pastur law, which characterizes the eigenvalue distribution of large random matrices, and related results from random matrix theory (RMT). However, training fundamentally alters these spectral properties, producing empirically observed ‘bulk+tail’ structured distributions that correlate strongly with generalization performance (Papyan, 2020; Martin and Mahoney, 2021). Existing theoretical frameworks fail to explain this transformation: while RMT describes initial conditions and stochastic differential equations (SDEs) can model SGD, current analyses focus on scalar parameters or low-rank models, failing to capture the full matrix-valued dynamics (Mandt et al., 2018). Most critically, no unified framework connects the microscopic stochastic dynamics of SGD to the macroscopic spectral evolution observed empirically.
In this paper, we bridge this gap by developing a continuous-time, matrix-valued SDE framework with carefully designed small-scale experiments that captures the full dynamics of singular value evolution under SGD. Our key contributions are:
  1. 1.
    We derive exact SDEs for individual singular values under isotropic SGD noise, connecting the microscopic parameter updates to macroscopic spectral dynamics. Under the assumption of negligible gradients, we show that squared singular values follow a Dyson Brownian motion with $\beta=1$, explaining the eigenvalue repulsion and spectral spreading.
  2. 2.
    We characterize the stationary spectral distribution in the non-negligible gradient regime using mean-field theory. We prove that the limiting distribution follows a gamma-type density with power-law tails, recovering the empirically observed ‘bulk+tail’ structure in trained networks and providing the first theoretical explanation for this empirical phenomenon.
  3. 3.
    Through experiments on transformer (Vaswani et al., 2017), vision transformer (Dosovitskiy et al., 2020), and MLP (Rumelhart et al., 1986) architectures, we demonstrate quantitative agreement between our SDE-based predictions and spectral evolution and propose an algorithm to forecast singular value dynamics from minimal gradient information.
We show that SGD’s stochastic noise acts like a “spectral sculptor”—initially spreading eigenvalues apart via repulsion, then concentrating them into beneficial empirically observed ‘bulk+tail’ structured patterns that enable generalization, connecting microscopic mini-batch randomness to macroscopic spectral evolution. Our work demonstrates how small-scale experimentation can unlock fundamental insights with implications for initialization strategies, optimization algorithm design, and understanding why deep learning works.
2 Related Works
Spectral Analysis of Neural Network Weights.
RMT establishes that at initialization, weight matrices follow Wigner’s semicircle law (Wigner, 1955) and the Marčenko–Pastur distribution (Marčenko and Pastur, 1967), with edge statistics governed by Tracy-Widom distributions (Tracy and Widom, 1996). Training induces pronounced deviations: singular-value spectra become highly anisotropic (Saxe et al., 2014), evolving ‘bulk+tail’ structured distributions linked to class structure (Papyan, 2020) and implicit regularization (Martin and Mahoney, 2021). Martin and Mahoney (Martin and Mahoney, 2021) identified 5+1 phases of spectral evolution and showed that batch size affects spectral properties, with smaller batches leading to stronger implicit self-regularization. Extensions to empirically observed ‘bulk+tail’ structured matrix ensembles (Auffinger et al., 2008) provide theoretical foundations past Gaussian universality classes.
SGD as Stochastic Dynamics.
SGD can be approximated by stochastic differential equations (SDEs), with constant-rate SGD behaving like an Ornstein–Uhlenbeck process (Mandt et al., 2017) and anisotropic noise structures enabling escape from sharp minima (Zhu et al., 2019). Weight updates have been mapped to Dyson Brownian motion (Aarts et al., 2025), explaining eigenvalue repulsion as a Coulomb-gas phenomenon. The mathematical foundation relies on Itô calculus for matrix functions (G. W. Stewart and Ji-guang Sun, 1990) and Fokker-Planck equations for interacting particle systems (Risken and Frank, 1996). The implicit regularization effects of gradient descent have been explored (Neyshabur et al., 2015).
3 Methodology
Our approach transitions from discrete microscopic SGD dynamics to continuous macroscopic spectral evolution. We produce notation for the training of neural networks via a spatiotemporal interpretation of the evolution of the weight matrices with stochasticity: $dW(x,t)=-\eta\frac{\partial\mathcal{L}}{\partial W(x,t)}dt+\sqrt{2\eta D_{W}(x,t)}\,d\mathcal{W}_{W}(x,t)$, where $\frac{\partial\mathcal{L}}{\partial W(x,t)}$ is the gradient of the loss, $\eta$ is the learning rate, $D$ is an effective diffusion constant described by $d\mathcal{W}_{W}$, and $d\mathcal{W}_{b}$ are independent matrix/vector-valued Wiener processes, capturing the stochastic dynamics of SGD. We split our analysis into two cases: negligible gradient in loss and non-negligible gradient in loss. For the first limit (negligible gradient in loss, or $\frac{\partial\mathcal{L}(W)}{\partial W}\approx 0$), we perform SVD/eigenvalue decomposition and use Ito Calculus (see Appendix 6.1) to arrive at the result of Theorem 3.1:
Theorem 3.1 (Stochastic Dynamics of Singular Values).
Let $W\in\mathbb{R}^{m\times n}$ evolve via stochastic gradient descent with noise. Then, the singular values $\sigma_{k}(W)$ follow the SDE:
$d\sigma_{k}(t)=\left[-\eta u_{k}^{T}(\nabla_{W}\mathcal{L})v_{k}+\eta D\left(\frac{m-n+1}{2\sigma_{k}}+\sum_{j\neq k}\frac{\sigma_{k}}{\sigma_{k}^{2}-\sigma_{j}^{2}}\right)\right]dt+\sqrt{2\eta D}d\beta_{k}(t)$
where $u_{k},v_{k}$ are the singular vectors and $D$ is the effective diffusion strength.
This theorem shows that under SGD noise, singular values behave like interacting particles that repel each other (the $\sum_{j\neq k}$ terms), explaining why the spectrum spreads out during training rather than collapsing. This SDE can then be mapped to Dyson-Brownian Motion processes (see Appendix 6.6 whose statistics are described by the Marčenko–Pastur (Lemma 6.9) and Tracy-Widom distributions respectively (Lemma 6.11).
After deriving these microscopic underpinnings, we may consider the second case (non-negligible gradient in loss, or $\frac{\partial\mathcal{L}(W)}{\partial W}\neq 0$) and transition into a macroscopic limit by considering the empirical spectral density distribution $\rho(\lambda,t)$ such that: $\rho(\lambda,t)=\frac{1}{r}\sum_{k=1}^{r}\delta(\lambda-\lambda_{k}(t))$ As $r\to\infty$, we assume that $\rho(\lambda,t)$ converges to a deterministic density function, normalized such that $\int\rho(\lambda,t)d\lambda=1$.
To study the dynamics of the squared singular values $\{\lambda_{j}\}$ then, we adopt a mean-field perspective by assuming the effective influence of the complex loss function $\mathcal{L}(W)$ is captured by a potential $\mathcal{L}_{MF}$ that depends only on this set. This postulates the form: $\mathcal{L}(W)\approx\mathcal{L}_{MF}(\{\lambda_{j}\})=\frac{c}{2}\sum_{j=1}^{r}(\lambda_{j}-\lambda^{*})^{2}$, where these eigenvalues $\lambda_{j}$ are “driven” to $\lambda^{*}$ as in Lemma 6.10.
Although deep neural network training objectives are generally nonconvex, it was shown in prior work (Ergen and Pilanci, 2025) that every stationary point of the original nonconvex problem coincides with the global optimum of a suitably defined subsampled convex program. As such, while the parameter-space landscape may admit many critical points, their spectral signatures at stationarity are governed by a convex variational principle. Consequently, modeling the large-$r$ limit of the empirical spectral density $\rho(\lambda,t)$ via a deterministic mean-field potential $\mathcal{L}_{MF}$ is fully justified as in Corollary 6.15. Evaluating this spectral density distribution, neglecting its highest order terms, and we derive Theorem 3.2:
Theorem 3.2 (Stationary Distribution of Singular Values).
Under the stationary mean-field approximation, the probability density function of the singular values follows a Gamma-type distribution:
$p_{\sigma}(\sigma)=2\frac{\left(\frac{\beta_{1}}{4\eta D}\right)^{\frac{m-n+3}{4}}}{\Gamma\left(\frac{m-n+3}{4}\right)}\sigma^{\frac{m-n+1}{2}}e^{-\left(\frac{\beta_{1}}{4\eta D}\right)\sigma^{2}}$
where $\beta_{1}$ is an effective noise constant representing the mean-field restoring force of the gradient, $D$ is the diffusion constant, and the weight matrix $W$ is $m\times n$.
This gamma-type distribution with power-law tails captures the “bulk+tail” structure observed empirically—most singular values cluster in a bulk region, while a few large values form the heavy tail that correlates with good generalization.
We present the proofs for both theorems in Appendix 6.1. Note that we treat Theorem 3.2 phenomenologically by fitting to within bounds beyond the Tracy-Widom distribution predictions (see Corollary 6.1). Thus, we take a “mean-field” approach to deconstructing this system. See Lemma 6.17 and Lemma 6.18 for the estimation of the effective diffusion and noise constants governing the stochastic dynamics.
4 Experiments
4.1 Experimental Setup
We validate our spectral‐SDE framework on three canonical architectures, namely a GPT-2 model; a Vision Transformer (ViT) (Dosovitskiy et al., 2020); and a MLP—all trained with SGD. This selection is motivated by extensive prior work showing ‘bulk+tail’ spectra arise across MLPs, CNNs, and transformers (Martin and Mahoney, 2021), and that different architectures exhibit distinct spectral biases and mode‐learning rates (Yao et al., 2022). We initialize all weights from architecture‐specific Gaussian priors, then train GPT-2 on Shakespeare text (Karpathy, 2015), and ViT/MLP on MNIST (LeCun et al., 2010) and CIFAR‐100 (Krizhevsky et al., 2009). More experimental details are in Appendix 8.
{comment}
4.2 Singular Value Evolution Simulation and Analysis
Figure 1: Singular‐value histograms at batches 0, 200, and 400, overlaid with the Marčenko–Pastur (MP) bulk law (red dashed) and the Tracy–Widom (TW) edge curve (green).
In Figure 1, at initialization the empirical spectrum adheres almost exactly to the MP prediction with no outliers beyond the TW edge, verifying random initialization assumptions. Our analysis begins by modeling the stochasticity of SGD with an isotropic noise term, resulting in a Langevin-type SDE. We acknowledge that this is an idealization; the true noise covariance of SGD is known to be anisotropic and parameter-dependent. However, this assumption provides a tractable starting point that allows us to establish a clear, analytical connection to the classical frameworks of Random Matrix Theory. This approach enables us to isolate and understand the fundamental repulsive dynamics that serve as a baseline for spectral evolution. We explicitly address the extension to the more realistic anisotropic case in our appendix (see Proposition 6.16), which we identify as a crucial direction for future work. At batch 200, the MP/TW fits begin to underpredict mass near the spectrum’s edge. This underprediction implies that growing correlations within the weight matrix diminish the effective size of the random matrix, consequently amplifying the dominance of edge statistics and fluctuations. Physically, this has the interpretation that we are still traversing either local minima or plateaus within the loss function landscape (producing weakly correlated learned features). At batch 400, this tail becomes increasingly pronounced: the bulk shifts rightward and a persistent shoulder of large singular values forms outside the MP support, showing that our matrix is becoming increasingly correlated. This is predicted to result from mostly gradient-derived information. Thus we understand the singular values beyond the TW threshold as highly correlated learned features. We empirically test our theory via Algorithm 1. The algorithm uses gradient loss and singular value information to use the dynamic equation previously described to predict singular values over time.
Figure 2: Predicted singular values (dashed) versus true.
Figure 3: Predicted heavy tails via Theorem 3.2.
In Figure 2, we track the top 8 singular values of a representative linear layer in a MLP over 800 training batches with CIFAR-10, plotting empirical trajectories against our bootstrap-drift predictions. The leading modes rise faster and the gradual increases of lower modes are reproduced by the prediction algorithm, with deviations starting around batch 300 as the spectrum begins to develop empirically observed ‘bulk+tail’ structure. This close alignment across all 8 modes up to the heavy‐tail regime demonstrates that our continuous‐time, matrix‐valued SDE framework accurately forecasts the full singular‐value dynamics from minimal gradient information. However, we hypothesize the anisotropic noise causes a bulk of the observed deviation for the singular values. This hypothesis is as follows: the larger singular values might experience greater effects from anisotropic noise due to preferential alignment of noise with their dominant singular vectors and potentially larger Hessian components (see Proposition 6.16). In Figure 3, the fits predict the qualitative shape well, but underpredict counts significantly, which we attribute to the aforementioned anisotropic noise hypothesis.
Figure 4: Mean variance for gaussian distributions
$\beta_{k}=\frac{1}{\sqrt{\eta\lambda_{k}}}\left(\frac{\partial\lambda_{k}}{\partial t}+\sqrt{\eta}\lambda_{k}\frac{\partial L}{\partial W}-\sum_{j\neq k}\frac{\lambda_{k}}{\lambda_{k}-\lambda_{j}}\right)$
4.3 Noise Analysis
As seen in the logarithmic plots above Figure 4 , the eigenvalue repulsive terms ($\sum_{j\neq k}\frac{\lambda_{k}}{\lambda_{k}-\lambda_{j}}$)) are dominant in the noise contribution, and the repulsive force term is strongly correlated with the other two forces ($\frac{\partial\lambda_{k}}{\partial t}+\sqrt{\eta}\lambda_{k}\frac{\partial L}{\partial W}$). If the noise from mini-batch sampling were truly random or isotropic (i.e., of equal magnitude in all directions), there would be no statistical basis for such a correlation. The fluctuations would be independent of the system’s internal dynamics. The existence of this correlation is therefore proof that the “random” kicks from SGD are systematically aligned with the geometry of the loss landscape and the current configuration of the eigenvalues. This suggests that the noise provides a sophisticated form of implicit regularization that goes beyond merely helping the model escape sharp minima. The noise actively reinforces the very dynamics that are beneficial for training. For instance, when two eigenvalues $\lambda_{i}$ and $\lambda_{j}$ get too close, the repulsive force between them increases. The observed correlation implies that the SGD noise will preferentially act in a direction that helps to separate them, effectively amplifying the natural repulsion.
4.4 Practical Applications
By understanding how singular values evolve during training, we can design spectral-aware initialization schemes that pre-configure empirically observed ‘bulk+tail’ structure bias to accelerate convergence, adaptive learning rate schedules that adjust $\eta/D$ ratios based on spectral concentration to balance exploration and convergence, and spectral-informed pruning that preserves these ‘bulk+tail’ structure contributors while aggressively compressing bulk parameters. Our findings also may enable novel optimizers that incorporate spectral information to treat parameters in large singular value directions differently than bulk directions, and spectral-based monitoring that tracks the transition from MP-like to empirically observed ‘bulk+tail’ structured spectra as an indicator of learning progress.
5 Conclusion
We develop a continuous-time, matrix-valued SDE framework connecting SGD’s microscopic dynamics to macroscopic spectral evolution, revealing that squared singular values follow Dyson Brownian motion and produce gamma-type distributions with power-law tails that explain the empirically observed ‘bulk+tail’ structure in trained networks. Through controlled experiments, we demonstrate quantitative agreement between our predictions and observed spectral evolution, with our forecasting algorithm accurately predicting singular value trajectories until empirically observed ‘bulk+tail’ structure emerges. While our current analysis assumes isotropic noise, future extensions to anisotropic SGD fluctuations could bridge the gap to real optimization dynamics and enable new preconditioning schemes.